# Characterization of the western Pictor A hotspot in the hard X-rays with NuSTAR

### v1

@@ 5 Conclusions
(i) We found that the NuSTAR data alone is best fit by a simple power law of photon index and a normalization at 1 keV of . We do not find any evidence for a spectral break or cutoff energy in the hard X-ray band. The difference in normalization between the recent 2021 observations and previous epochs of both Chandra and NuSTAR data is suggests variability within the western hot spot.
@@
The emerging picture of the hot spot is of multiple localized, highly magnetized, pc-scale substructures emitting X-ray predominantly via synchrotron radiation. This paradigm is capable of explaining the rapid X-ray variability and the difference in spectral index between the radio and X-rays. Magnetic reconnection models are capable of reproducing the necessary in situ acceleration in smaller regions.
@@ Appendix A NuSTAR background models
The nuskybgd package uses XSPEC333https://heasarc.gsfc.nasa.gov/xanadu/xspec (Arnaud, 1996) to fit the spectrum of the entire background across the NuSTAR detector modules, taking into account each component and any spatial variations present. nuskybgd groups the input spectrum to a minimum of 30 counts per bin and calculates the parameter normalizations using the statistic. Within XSPEC, the background model is defined using a combination of four different models. apbgd and fxapbgd model the ‘aperture’ and ‘focused aperture’ or ‘fCXB’ components and are both defined as cutoff power-law components (cutoffpl) with a photon index of and a high-energy cutoff of keV. The line_bgd component models the fluorescense and activation line elements of the ‘internal’ background component and the Solar thermal emission of the ‘scattered’ background. It is comprised of 29 different Lorentzian (lorentz) line profiles and a thermal plasma component (apec) with temperature . The final element is the particle_bgd component which models the featureless continuum of the ‘internal’ background as a broken power-law with photon indices of and a break enery of keV.
@@ Appendix B Results for individual Chandra epochs
CH-1 | PL* | - | - | - | 0.00 | 0.416 | 62.4/69 = 0.904 | |||
Bkn. PL | - | -0.05 | 0.370 | 61.1/67 = 0.912 | ||||||
Dbl. Bkn. PL | -0.34 | 0.190 | 63.6/65 = 0.979 | |||||||
CH-2 | PL | - | - | - | -0.50 | 0.156 | 116.1/160 = 0.726 | |||
Bkn. PL* | - | 0.00 | 0.492 | 112.1/158 = 0.709 | ||||||
Dbl. Bkn. PL | -0.19 | 0.318 | 110.6/156 = 0.709 | |||||||
CH-3 | PL* | - | - | - | 0.00 | 0.348 | 97.2/124 = 0.784 | |||
Bkn. PL | - | -0.03 | 0.325 | 97.7/122 = 0.801 | ||||||
Dbl. Bkn. PL | -0.09 | 0.283 | 97.4/120 = 0.812 | |||||||
CH-4 | PL* | - | - | - | 0.00 | 0.373 | 20.7/24 = 0.862 | |||
Bkn. PL | - | -0.02 | 0.356 | 19.9/22 = 0.904 | ||||||
Dbl. Bkn. PL | -0.18 | 0.246 | 21.0/20 = 1.051 | |||||||
CH-5 | PL* | - | - | - | 0.00 | 0.391 | 38.9/75 = 0.519 | |||
Bkn. PL | - | 0.00 | 0.365 | 39.6/73 = 0.543 | ||||||
Dbl. Bkn. PL | -0.24 | 0.225 | 39.5/71 = 0.556 | |||||||
CH-6 | PL | - | - | - | -0.08 | 0.304 | 43.1/63 = 0.684 | |||
Bkn. PL* | - | 0.00 | 0.365 | 42.7/61 = 0.700 | ||||||
Dbl. Bkn. PL | -0.35 | 0.163 | 41.9/59 = 0.711 | |||||||
CH-7 | Exp. Cutoff PL | - | - | -0.02 | 0.294 | 59.7/70 = 0.852 | ||||
Bkn. PL* | - | 0.00 | 0.308 | 57.9/69 = 0.839 | ||||||
Dbl. Bkn. PL | -0.24 | 0.177 | 57.2/67 = 0.854 |
Appendix C Image deconvolution with LIRA
We use a statistical tool called Low Count Image Reconstruction and Analysis (LIRA, Esch et al., 2004; Stein et al., 2015) to deconvolve images. Although LIRA was originally designed to detect the statistical significance of structures in low-count high-energy observations, it also has been used to deconvolve sky images (e.g., Kashyap et al., 2017; Reddy et al., 2021b, 2023). We refer the reader to Stein et al. (2015) for a full statistical treatment. Briefly, LIRA models the brightness distribution of an observation using two model images. One is a baseline with known sources of emission (e.g., core and background), which is specified by the user, and an added model, which is to be inferred containing any emission over the baseline, required to explain the observation. We can reconstruct the observation by convolving their sum with the PSF. LIRA performs Bayesian analysis using a flexible multi-scale representation of the observation to infer the posterior distribution of the added model and samples it with a Markov Chain Monte Carlo (MCMC) sampler to generate a series of images. These images represent the observation’s brightness distribution after subtracting the baseline. With nuskybgd-generated background as the baseline, we generate 2000 images after discarding the first 1000 images as burn-in and average them to produce a deconvolved image.
Due to the spatial dependence of NuSTAR’s PSF, using a single PSF to deconvolve both the core and the hot spot, which are separated by 4’, may lead to artifacts in the output images. Hence, we separately deconvolve the core and the hot spot regions. Furthermore, to determine the energy dependence of the structure in the jet and the hot spot, we divide the image into two energy ranges of 3-12 keV and 12-78 keV, as shown in the top and bottom rows of figure A2. The left and right two panels in each row show core and WHS images, respectively. We set the core image size to 128x128 and WHS to 64x64, following the input requirements of LIRA.
The top and bottom rows of panels in figure A2 show the sky and LIRA-deconvolved images of the core and the WHS in the energy ranges 3-12 keV and 12-78 keV, respectively. We find no evidence for an inner jet in the core’s deconvolution in both the energy ranges (panels b and f), indicating its emission is within the error limits of the core’s PSF and background. This absence of emission is expected in the sense that the flux from the inner jet is close to the detection limit of NuSTAR. The 0.5-7 keV energy flux of the brightest feature in the inner jet is erg cm s, while the 3-sensitivity of NuSTAR with a 1 Ms exposure444Measured in the 6-10 keV band (about three times larger than our combined 300 ks exposure) is erg cm s (Harrison et al., 2013). Finally, because we also find no evidence for a jet beyond 12 keV (fig. A2(e)), we use the estimated 10-30 keV sensitivity of erg cm s as an upper limit for each knot. Hence, we take erg cm s as an upper limit on the inner jet’s 10-30 keV flux.

### v2

@@ 5 Conclusions
(i) We found that the NuSTAR data alone is best fit by a simple power law of photon index and a normalization at 1 keV of . We do not find any evidence for a spectral break or cutoff energy in the hard X-ray band. The difference in normalization between the recent 2021 observations and previous epochs of both Chandra and NuSTAR data suggests variability within the western hotspot.
@@
The emerging picture of the hotspot is of multiple localized, highly magnetized, pc-scale substructures emitting X-ray predominantly via synchrotron radiation in agreement with previous research (e.g., Tingay et al., 2008; Isobe et al., 2020). This paradigm is capable of explaining the rapid X-ray variability and the difference in spectral index between the radio and X-rays. Magnetic reconnection models are a possible explanation for in situ acceleration in smaller regions and can produce electrons at the required TeV energies (e.g. Werner et al., 2016; Bodo et al., 2021).
@@ Appendix A NuSTAR background models
The nuskybgd package uses XSPEC444https://heasarc.gsfc.nasa.gov/xanadu/xspec (Arnaud, 1996) to fit the spectrum of the entire background across the NuSTAR detector modules, taking into account each component and any spatial variations present. nuskybgd groups the input spectrum to a minimum of 30 counts per bin and calculates the parameter normalizations using the statistic. Within XSPEC, the background model is defined using a combination of four different models. apbgd and fxapbgd model the ‘aperture’ and ‘focused aperture’ or ‘fCXB’ components and are both defined as cutoff power-law components (cutoffpl) with a photon index of and a high-energy cutoff of keV. The line_bgd component models the fluorescense and activation line elements of the ‘internal’ background component and the Solar thermal emission of the ‘scattered’ background. It is comprised of 29 different Lorentzian (lorentz) line profiles and a thermal plasma component (apec) with temperature . The final element is the particle_bgd component which models the featureless continuum of the ‘internal’ background as a broken power-law with photon indices of and a break enery of keV.